# Training Gradient Boosted Decision Trees on Tabular Data Containing Label Noise for Classification Tasks

### v1

@@ 7 Conclusion
This work explores the effects of label noise on gradient boosted decision trees (GBDTs) and compares different methods of label noise detection and correction. The primary objectives, as outlined in Section 1.1, are to provide an overview of the label noise research landscape, implement a GBDT which improves upon other GBDTs trained on noisy, tabular data and evaluate its performance on benchmark datasets.
The literature review provided a structured overview of label noise research, categorizing it by their general approach. Then it highlighted the most recent advancements in the field, namely the training of deep neural networks (DNNs) on noisy image and text data. Training GBDTs on tabular data was then identified as a notable gap in research. Stemming from this observation, the primary research questions were formulated. These encompass studying the impact of label noise on GBDTs, adapting certain methodologies from DNN research to GBDTs, and subsequently evaluating their effectiveness on benchmark datasets, with a comparative analysis against other state-of-the-art research.
This was achieved by first gathering data, preprocessing it, training GBDTs at various noise rates and types and finally interpreting the results both visually and numerically. Four benchmark datasets, namely the Breast Cancer, Covertype, Adult, and Dry Bean dataset, were used to evaluate the implemented methods. The first three datasets served as a basis for comparison with state-of-the-art work, while the Dry Bean and Covertype datasets helped explore the behavior of GBDTs under label noise conditions.
The experiments revealed that GBDTs are robust to label noise in the early stages of training, especially to symmetric noise, and can achieve good performance by utilizing early stopping. In the scope of this work, methods from deep learning were adapted and both AUM and LRT achieved state-of-the-art noise detection performance on the Adult dataset. The percentage of instances marked as noisy by AUM and LRT can be utilized to approximate the level of noise present in the data, provided that the dataset is not excessively imbalanced or contains too many outliers. But it is still advisable to treat fewer instances than the estimated amount of noise present in the data.
@@
Inconsistency in performance of the methods across various datasets implies that the success of the implemented label noise detection and correction methods could depend on factors such as the dataset, noise rate, noise type, and the chosen correction technique. Alternatively, these favorable results might be attributed to random chance rather than the inherent effectiveness of the methods. Conducting repeated experiments and employing cross-validation could resolve these ambiguities.
@@
Future research could explore other relabeling methods, beyond using the most confident or frequent predictions from recent epochs. And because high noise detection performance didn’t always translate in a high increase in classification accuracy, further measures like error correction may be needed to effectively make use of the noise detection methods. Since classification performance decreased when removing or relabeling noisy candidates on imbalanced datasets, future research could also investigate methods to reliably detect noisy instances on data with high class imbalance using GBDTs. Until now, GBDTs generally outperform DNNs on tabular data, but if that were to change it would be interesting to investigate whether label noise treatment methods from image and text classification are effective on tabular data as well.
In summary, this work conducted a comprehensive analysis of the impact of label noise on GBDTs, offering practical guidance. Two methods for dealing with label noise from the area of deep learning were successfully adapted to work with GBDTs. Building on the work of Ponti et al. (2022), their noise detection algorithm was combined with automatic relabeling of noisy candidate instances. These modified GBDTs were tested against benchmark datasets, and their performance was compared to current state-of-the-art approaches. The findings here provide a basis for continued exploration and improvement in handling label noise, particularly in the context of tabular data.
8 CRediT authorship contribution statement
@@
Anita Eisenbürger: Conceptualization, Data Curation, Formal analysis, Investigation, Methodology, Visualization, Writing – original draft, Writing – review & editing. Daniel Otten: Supervision, Writing – review & editing. Frank Hopfgartner: Supervision, Writing – review & editing. Anselm Hudde: Writing – review & editing.

### v2

@@ 8 Conclusion
This work examines the impact of label noise on Gradient Boosted Decision Trees (GBDTs) and evaluates several label noise detection and correction methods tailored for tabular data—an underexplored area in label noise research. We adapted two noise detection techniques from deep learning for use with GBDTs and further developed two methods specifically for this study: the Gradients method, inspired by weight clipping in Adaboost and the small loss trick, and an extension of ConfCorr to support relabeling.
Our experiments demonstrate that GBDTs exhibit natural robustness to label noise in the early training stages, particularly in cases of symmetric noise, where early stopping proved effective in preserving test accuracy. Both AUM and LRT achieved state-of-the-art noise detection accuracy on the Adult dataset, with the percentage of instances marked as noisy by these methods providing a useful estimate of the dataset’s noise level. In addition, ConfCorr achieved high classification precision on the Adult dataset, while Gradients outperformed other methods in terms of classification precision and recall on the Covertype and Breast Cancer datasets, respectively.
@@
The inconsistency in method performance across different datasets indicates that optimal noise detection and correction strategies may depend on factors such as dataset characteristics, noise rate, and noise type. This variability suggests that favorable results may, in some cases, stem from dataset-specific factors rather than inherent method effectiveness. Repeated experiments and cross-validation could provide greater clarity and reliability in assessing method performance.
@@
Future research could explore alternative relabeling strategies beyond relying on the most confident or frequent predictions from recent epochs. Since high noise detection accuracy does not always lead to improved classification performance, further corrective measures, such as error correction, may enhance the utility of noise detection methods. Additionally, given that classification performance often decreased when removing or relabeling noisy instances in imbalanced datasets, future studies could develop noise detection techniques better suited to class-imbalanced data in GBDTs.
As GBDTs continue to generally outperform Deep Neural Networks (DNNs) on tabular data, this work contributes essential insights into label noise handling for these models. Should DNNs eventually surpass GBDTs in tabular tasks, it would be valuable to investigate the applicability of noise handling techniques from image and text domains for tabular data. This study serves as a foundational step in advancing robust GBDT training under noisy conditions, laying the groundwork for continued exploration in this critical area of machine learning.
9 CRediT Authorship Contribution Statement and AI Disclosure
@@
Generative AI tools, such as ChatGPT, were used under human oversight solely to improve the readability and language of this manuscript. No generative AI was used for data analysis, content generation, or any other aspect of the research process. Anita Eisenbürger: Conceptualization, Data Curation, Formal analysis, Investigation, Methodology, Visualization, Writing – original draft, Writing – review & editing. Daniel Otten: Supervision, Writing – review & editing. Frank Hopfgartner: Supervision, Writing – review & editing. Anselm Hudde: Writing – review & editing.